find the gradient of the line that goes through (-3,-5) and (-1,-7). and hence, find the equation of this traight line.
Answer by Cromlix(4381)

Using equation: y2 - y1/x2 - x1
-7 -(-5)/-1 - (-3)
-2/2 = -1 (gradient)
Using equation:
y - b = m(x - a)
a = -5 b = -3 m = -1
y - (-3)= -1(x - (-5))
y + 3 = -1(x + 5)
y + 3 = -x - 5
y = -x - 5 - 3
y = -x - 8
Hope this helps.
:-)